# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

### v2

@@ Appendix D Illustrative Case Study Example
We consider a benchmark of verbal reasoning puzzles drawn from a long-running radio show wu2025phd. The original authors distinguish two settings: (i) on-air puzzles, which are short, pattern-based word questions intended to be solved live in a few seconds by callers during the broadcast, and (ii) off-air weekly challenges, which are substantially harder puzzles that listeners may think about for days before submitting an answer. In the released benchmark, GPT-4o achieves around 65% accuracy on the easier on-air subset, but only about 6% on the more challenging off-air subset.
@@ D.2 MMLU-Pro: Choosing the number of answer options
To illustrate how description-based prediction can support design-time decisions, we consider a multitask multiple-choice benchmark in the style of MMLU-Pro (wang2024mmlu), and focus on a single design configuration: the number of answer options per question. We imagine a benchmark designer who has not yet built the dataset and is deciding between 5, 10, 15, 20, or 25 options. For each configuration, the designer writes a natural language description (same questions and domains, different number of options) and queries our predictor, which estimates GPT–4o (hurst2024gpt) accuracy from text alone, without access to any items.
Figure 7 summarize the results. Using only these textual descriptions, our predictor (GPT–5) estimates GPT–4o accuracies of 86%, 70%, 65%, 62%, and 60% for 5, 10, 15, 20, and 25 options, respectively, while the reported ground truth accuracy at 10 options is 72.6% from the original paper (wang2024mmlu). The predictions are monotone in the expected direction and reasonably calibrated at the configuration used in practice.

### v3

@@ Appendix D Illustrative Case Study Example
We consider a benchmark of verbal reasoning puzzles drawn from a long-running radio show (Wu et al., 2025). The original authors distinguish two settings: (i) on-air puzzles, which are short, pattern-based word questions intended to be solved live in a few seconds by callers during the broadcast, and (ii) off-air weekly challenges, which are substantially harder puzzles that listeners may think about for days before submitting an answer. In the released benchmark, GPT-4o achieves around 65% accuracy on the easier on-air subset, but only about 6% on the more challenging off-air subset.
@@ D.2 MMLU-Pro: Choosing the number of answer options
To illustrate how description-based prediction can support design-time decisions, we consider a multitask multiple-choice benchmark in the style of MMLU-Pro (Wang et al., 2024), and focus on a single design configuration: the number of answer options per question. We imagine a benchmark designer who has not yet built the dataset and is deciding between 5, 10, 15, 20, or 25 options. For each configuration, the designer writes a natural language description (same questions and domains, different number of options) and queries our predictor, which estimates GPT–4o (Hurst et al., 2024) accuracy from text alone, without access to any items.
Figure 7 summarize the results. Using only these textual descriptions, our predictor (GPT–5) estimates GPT–4o accuracies of 86%, 70%, 65%, 62%, and 60% for 5, 10, 15, 20, and 25 options, respectively, while the reported ground truth accuracy at 10 options is 72.6% from the original paper (Wang et al., 2024). The predictions are monotone in the expected direction and reasonably calibrated at the configuration used in practice.